OpenStudy (tia_sweetheart17):

write each product as a rational expression reduced to lowest terms. 42/7 times 35/49

OpenStudy (campbell_st):

simplify each fraction 1st by cancelling common factors \[42/7 = (6 \times 7)/ 7 = 6\] \[35/49 = (5 \times 7)/(7\times7) = 5/7\] \[42/7 \times 35/49 = 6 \times 5/7 = 30/7\]

OpenStudy (campbell_st):

lol... luis I think its 4 and 2/7.... looks like a typo